# Camouflaged Object Tracking: A Benchmark

### v1

@@ 7.2 The Co-occurrence Distributions of the 11 Attributes in COTD
8 Additional Results
8.1 Overall Performance
The normalized precision curve of all trackers is shown in Fig.12. Notably, HIPTrack-MLS achieves superior performance compared to all other trackers, with a nomalized precision of 0.824, which is 0.8% higher than its baseline, HIPTrack, which scores 0.816.
8.2 Attribute-Based Performance
From Fig.14, we observe that HIPTrack-MLS ranks first in the majority of the 12 attributes, including BC, aligning with the outcomes from the comprehensive test set evaluation.
8.3 More Qualitative Results
Table 4: Overall tracking results of the trackers on Non-aquatic and Aquatic,The first-, second- and third-place trackers are labeled with red, blue and green colors respectively.
@@ 8.3 More Qualitative Results
Aba-ViTrack[43] ICCV 2023 56.34 33.48 64.80 67.91 40.04 81.56
AQATrack[75] CVPR 2024 61.74 41.73 71.53 75.72 56.22 88.82
ARTrack[68] CVPR 2023 61.18 40.43 69.80 75.76 58.42 88.83
CSWinTT[65] CVPR 2022 56.67 32.04 64.04 68.56 43.10 81.76
DCPT[92] ICRA 2024 60.27 38.71 68.32 73.41 52.04 86.75
DropTrack[69] CVPR 2023 63.03 42.61 73.08 75.03 56.14 87.65
GRM[27] CVPR 2023 59.59 38.29 69.06 73.29 53.99 86.56
@@
HIPTrack[5] CVPR 2024 63.95 43.69 74.22 75.42 56.42 88.01
JointNLT[90] CVPR 2023 44.41 19.94 50.46 58.99 27.45 70.95
MAT[87] CVPR 2023 56.86 33.47 64.35 64.92 37.25 78.03
ROMTrack[7] ICCV 2023 61.16 40.59 71.77 74.90 58.47 87.46
SeqTrack [11] CVPR 2023 59.23 40.63 70.22 72.29 53.03 86.24
@@
In our research, we meticulously conduct a comprehensive experiment to evaluate the performance of multiple state-of-the-art trackers when handling camouflaged objects, with the results systematically presented in Table 4. We strategically categorize the COTD dataset into two superclasses—Aquatic and Non-aquatic—to thoroughly test the capabilities of 20 state-of-the-art trackers, alongside our tracker, HIPTrack-MLS. The experiment hinges on three pivotal metrics: AUC, Precision, and Normalized Precision, to provide a robust assessment of each tracker’s performance. The experimental results demonstrate that our improved tracker HIPTrack-MLS excels in all evaluated metrics, significantly outperforms its baseline, HIPTrack. Specifically, in the domain of non-aquatic object tracking, HIPTrack-MLS achieves an AUC that is 0.58% higher than HIPTrack, with an increase from 63.95% to 64.53%. Additionally, HIPTrack-MLS also leads with a precision of 44.52%, surpassing the second-place HIPTrack, which has a precision of 43.69%, by 0.83%. While there is a minor 0.12% reduction in normalized precision, from 74.22% to 74.10%, this decrement is negligible in the context of real-world applications. When it comes to the aquatic objects, HIPTrack-MLS with a AUC of 77.21%, which is 1.79% higher than its baseline, HIPTrack, which scores 75.42%. For the precision and normalized precision, HIPTrack-MLS shows improvements of 1.08% and 1.48%, respectively, over HIPTrack. These results indicate that our tracker performs exceptionally well in the tracking of camouflaged objects. Additionally, we observe performance variances among other trackers in both objects. Some trackers, such as AQATrack and HIPTrack, show higher precision and normalized precision in aquatic objects, while others, like HiFT and SiamGAT, exhibit relatively lower AUC in non-aquatic objects. This observation potentially indicates limitations in their generalization capabilities and underscores the importance of further research in this area.
@@
To make sense of the advantage of the propose method, we explore the allocation of attention in the task of Camouflaged Object Tracking and compare the performance of the improved HIPTrack-MLS model with the original HIPTrack model. Fig. 13 intuitively displays three types of visual representations: the original image, the attention map of the original HIPTrack model, and the attention map of the HIPTrack-MLS model. The original images in the first row demonstrate the camouflage of the object in a complex background, which poses a significant challenge to tracking algorithms. The high degree of visual integration between the object and the background increases the difficulty of tracking. The second row shows the attention maps generated by the original HIPTrack model. These maps reveal some limitations of the original model when handling camouflaged objects, especially when the similarity between the object and the background is high, the original model may struggle to maintain precise tracking of the object. The third row, however, showcases the attention maps of the HIPTrack-MLS model, which clearly highlight the target object. Even when the object is camouflaged, the model can accurately identify and focus attention on the key features of the object. This improved attention mechanism significantly enhances tracking accuracy, especially in complex environments.
Figure 14: Performance of 12 attributes on COTD (shown using a success plot). Best viewed by zooming in.

### v2

@@ V-B Evaluation Results
Evaluation on different object categories: In our research, we meticulously conduct a comprehensive experiment to evaluate the performance of multiple state-of-the-art trackers when handling camouflaged objects, with the results systematically presented in Table IV. We strategically categorize the COTD dataset into two superclasses—Aquatic and Non-aquatic—to thoroughly test the capabilities of 20 state-of-the-art trackers, alongside our tracker, HIPTrack-MLS. The experiment hinges on three pivotal metrics: AUC, Precision, and Normalized Precision, to provide a robust assessment of each tracker’s performance. The experimental results demonstrate that our improved tracker HIPTrack-MLS excels in all evaluated metrics, significantly outperforms its baseline, HIPTrack. Specifically, in the domain of non-aquatic object tracking, HIPTrack-MLS achieves an AUC that is 0.58% higher than HIPTrack, with an increase from 63.95% to 64.53%. Additionally, HIPTrack-MLS also leads with a precision of 44.52%, surpassing the second-place HIPTrack, which has a precision of 43.69%, by 0.83%. While there is a minor 0.12% reduction in normalized precision, from 74.22% to 74.10%, this decrement is negligible in the context of real-world applications. When it comes to the aquatic objects, HIPTrack-MLS with a AUC of 77.21%, which is 1.79% higher than its baseline, HIPTrack, which scores 75.42%. For the precision and normalized precision, HIPTrack-MLS shows improvements of 1.08% and 1.48%, respectively, over HIPTrack. These results indicate that our tracker performs exceptionally well in the tracking of camouflaged objects. Additionally, we observe performance variances among other trackers in both objects. Some trackers, such as AQATrack and HIPTrack, show higher precision and normalized precision in aquatic objects, while others, like HiFT and SiamGAT, exhibit relatively lower AUC in non-aquatic objects. This observation potentially indicates limitations in their generalization capabilities and underscores the importance of further research in this area.
TABLE IV: Overall tracking results of the trackers on Non-aquatic and Aquatic,The first-, second- and third-place trackers are labeled with red, blue and green colors respectively.
@@
STARK[87] ICCV 2021 57.93 35.01 66.42 69.80 42.16 84.58
@@
Aba-ViTrack[96] ICCV 2023 56.34 33.48 64.80 67.91 40.04 81.56
ROMTrack[72] ICCV 2023 61.16 40.59 71.77 74.90 58.47 87.46
ARTrack[74] CVPR 2023 61.18 40.43 69.80 75.76 58.42 88.83
DropTrack[73] CVPR 2023 63.03 42.61 73.08 75.03 56.14 87.65
GRM[76] CVPR 2023 59.59 38.29 69.06 73.29 53.99 86.56
JointNLT[91] CVPR 2023 44.41 19.94 50.46 58.99 27.45 70.95
MAT[89] CVPR 2023 56.86 33.47 64.35 64.92 37.25 78.03
SeqTrack [75] CVPR 2023 59.23 40.63 70.22 72.29 53.03 86.24
@@
DCPT[86] ICRA 2024 60.27 38.71 68.32 73.41 52.04 86.75
AQATrack[84] CVPR 2024 61.74 41.73 71.53 75.72 56.22 88.82
HIPTrack[71] CVPR 2024 63.95 43.69 74.22 75.42 56.42 88.01
@@
TABLE V: Efficiency comparison evaluated on COTD.
HIPTrack-MLS HIPTrack ProContEXT ZoomTrack DCPT
FPS 15.67 23.46 10.87 15.62 11.25
@@
Attention mechanism analysis: To make sense of the advantage of the propose method, we explore the allocation of attention in the task of Camouflaged Object Tracking and compare the performance of the improved HIPTrack-MLS model with the original HIPTrack model. Fig. 12 intuitively displays three types of visual representations: the original image, the attention map of the original HIPTrack model, and the attention map of the HIPTrack-MLS model. The original images in the first row demonstrate the camouflage of the object in a complex background, which poses a significant challenge to tracking algorithms. The high degree of visual integration between the object and the background increases the difficulty of tracking. The second row shows the attention maps generated by the original HIPTrack model. These maps reveal some limitations of the original model when handling camouflaged objects, especially when the similarity between the object and the background is high, the original model may struggle to maintain precise tracking of the object. The third row, however, showcases the attention maps of the HIPTrack-MLS model, which clearly highlight the target object. Even when the object is camouflaged, the model can accurately identify and focus attention on the key features of the object. This improved attention mechanism significantly enhances tracking accuracy, especially in complex environments.
TABLE VI: Performance comparison of HIPTrack-MLS and its baseline HIPTrack on generic and our benchmarks.
Tracker LaSOT TrackingNet COTD
Prec. Succ. Prec. Succ. Prec. Succ.
HIPTrack 0.795 0.727 0.838 0.845 0.515 0.680
HIPTrack-MLS 0.793 0.724 0.837 0.845 0.525 0.692
Efficiency analysis: To evaluate the computational efficiency of our proposed method, we conducted runtime analysis on the COTD benchmark dataset. Table Table V presents the tracking speed in frames per second (FPS) for HIPTrack-MLS and other state-of-the-art trackers. The results show that HIPTrack-MLS achieves a tracking speed of 15.67 FPS, which is comparable to ZoomTrack (15.62 FPS) and faster than ProContEXT (10.87 FPS) and DCPT (11.25 FPS). While the original HIPTrack demonstrates higher efficiency at 23.46 FPS, our multi-level feature sensing architecture introduces additional computational overhead due to the feature fusion and dynamic weighting mechanisms. However, this trade-off between speed and accuracy is justified by the significant performance improvements in tracking camouflaged objects, as demonstrated in our previous experiments.
Performance on generic datasets: To validate whether HIPTrack-MLS’s improvements are specifically beneficial for camouflaged object tracking or represent a general enhancement in tracking capability, we conduct comprehensive evaluations on standard tracking benchmarks including LaSOT and TrackingNet, alongside our COTD dataset. The results are presented in Table VI. On the LaSOT dataset, HIPTrack-MLS achieves comparable performance to the baseline HIPTrack, with precision and success scores of 0.793 and 0.724, respectively, showing a slight marginal difference compared to HIPTrack’s 0.795 and 0.727. Similarly, on TrackingNet, HIPTrack-MLS demonstrates minimal improvement with precision and success scores of 0.837 and 0.845, compared to HIPTrack’s 0.838 and 0.845. However, on the COTD dataset, HIPTrack-MLS shows a significant performance gain, improving the success rate from 0.680 to 0.692 and precision from 0.515 to 0.525. This notable improvement on COTD, coupled with maintained performance on general tracking benchmarks, suggests that our multi-level feature sensing architecture specifically addresses the challenges of camouflaged object tracking without compromising general tracking capability.
The experimental results support our hypothesis that the multi-level feature sensing mechanism is particularly effective for camouflaged objects, where fine-grained feature discrimination at different scales is crucial for maintaining tracking accuracy. The maintained performance on general tracking benchmarks also indicates that our modifications do not introduce any degradation in general tracking scenarios, making HIPTrack-MLS a versatile tracking solution.
Qualitative evaluation: In Fig. 11, we demonstrate qualitative comparisons of our method with 9 top trackers, including HIPTrack [71], DropTrack [73], DCPT [86], AQATrack [84], ARTrack [74], GRM [76], ROMTrack [72], SeqTrack [75], and ZoomTrack [88]. These examples face challenges such as Illumination Variation, Deformation, Rotation, Aspect Ratio Change, and Partial Occlusion. Specifically, in the fish-4 sequence, only HIPTrack-MLS consistently tracks the camouflaged fish. In mantis-2, while some trackers manage to track the mantis, HIPTrack-MLS consistently demonstrates superior visual performance. In person-9, due to background clutter and multiple object interference, only HIPTrack-MLS, DropTrack, ARTrack, and SeqTrack can track the target person. In phasmatodea-10, many trackers are affected by the background, leading to inaccurate tracking of the camouflaged phasmatodea. These qualitative results validate the superiority of our method for tracking camouflaged objects.
V-C Ablation study
Impact of the importance coefficient in Multi-Level Feature Sensing. To examine the effect of the weight coefficient $\gamma$ on balancing the contributions of the aggregated multi-level features and $H_{n}^{\rightarrow}$ in HIPTrack-MLS, we evaluated HIPTrack-MLS on the COTD dataset with a range of $\gamma$ settings. Specifically, $\gamma$ ranges from 0.1 to 1.0 with increments of 0.1. As $\gamma$ increases, the influence of the aggregated feature becomes more pronounced, while the reliance on $H_{n}^{\rightarrow}$ decreases. Table VII details HIPTrack-MLS’s PRC and AUC on the COTD dataset for various $\gamma$ values. The results indicate that the optimal PRC and AUC, which are 52.5% and 69.2% respectively, are achieved when $\gamma$ is set to 0.1. This result indicates that an appropriate choice of $\gamma$ enhances tracking performance, confirming the effectiveness of the proposed method. Consequently, we adopted $\gamma=0.1$ as the default setting for HIPTrack-MLS.
TABLE VII: Illustrates the impact of the importance coefficient of aggregated feature on HIPTrack-MLS’s PRC (%) and AUC (%) on COTD. Red, blue, and green represent the first, second, and third place, respectively.
$\gamma$ 1.0 0.9 0.8 0.7 0.6 0.5 0.4 0.3 0.2 0.1
PRC (%) 47.0 48.1 49.1 49.9 50.4 51.1 51.7 52.1 52.2 52.5
AUC (%) 66.7 67.4 68.0 69.1 68.1 68.2 68.7 68.8 68.3 69.2
VI Conclusion